# Improving Robustness and Reliability in Medical Image Classification with Latent-Guided Diffusion and Nested-Ensembles

### v4

@@ Abstract
Ensemble deep learning has been shown to achieve high predictive accuracy and uncertainty estimation in a wide variety of medical imaging contexts. However, perturbations in the input images at test time (e.g. noise, domain shifts) can still lead to significant performance degradation, posing challenges for trustworthy clinical deployment. In order to address this, we propose LaDiNE, a novel and robust probabilistic method that is capable of inferring informative and invariant latent variables from the input images. These latent variables are then used to recover the robust predictive distribution without relying on a predefined functional-form. This results in improved (i) generalization capabilities and (ii) calibration of prediction confidence. Extensive experiments were performed on the task of disease classification based on the Tuberculosis chest X-ray and the ISIC Melanoma skin cancer datasets. Here the performance of LaDiNE was analysed under a range of challenging covariate shift conditions, where training was based on “clean” images, and unseen noisy inputs and adversarial perturbations were presented at test time. Results show that LaDiNE outperforms existing state-of-the-art baseline methods in terms of accuracy and confidence calibration. This increases the feasibility of deploying reliable medical machine learning models in real clinical settings, where accurate and trustworthy predictions are crucial for patient care and clinical decision support.
Index Terms:
@@ I Introduction
In the rapidly evolving domain of medical imaging analysis, deep learning has led to enormous advances in many clinical domains of interest [1, 2, 3, 4, 5, 6, 7, 8, 9], notably in tasks such as detection of diabetic retinopathy in eye fundus images [2], classification of skin cancer [1], and identification of cancerous regions in mammograms [7]. Despite the fact that recent methods have achieved unprecedented success in controlled experimental settings, their fundamental building blocks – deep neural networks (DNNs) – are known to be sensitive to slight distribution changes and vulnerable to attacks [10]. As a result, their application to real-world clinical contexts often results in significant performance degradation, including inaccurate predictions and poorly calibrated confidence estimates. This results in mistrust by clinicians in deploying them in real clinical settings [11, 12].
@@
Data augmentation is a widely used tool for improving the generalization of DNNs. In the field of medical imaging, however, where datasets are typically smaller than natural image datasets especially for rare diseases [13], conventional data augmentation strategies may not always be suitable and can even lead to degradation in the performance of DNNs [14]. Designing suitable augmentation strategies from small medical imaging datasets can be challenging, requiring careful design choices in order incorporate suitable inductive biases that lead to robust and generalizable DNN architectures and learning algorithms.
@@
Ensemble methods are popular choices to enhance generalization, as they combine the predictions of multiple models, effectively reducing variance and mitigating overfitting [15]. By leveraging the strengths of diverse models, ensemble techniques such as bagging, boosting, and stacking can achieve better and more robust predictive performance over a single model [16]. In addition, deep ensemble methods have been shown to improve both predictive performance and the quality of uncertainty estimates by training multiple DNNs independently and averaging their predictions [17]. Other frameworks involve developing algorithms to select better ensemble members and distribute input-dependent weights to each member, aiming to reduce the effect of weak members and providing performance gains [18, 19]. However, despite these advances, ensemble methods often rely on restricted simple component distribution assumptions (e.g., Gaussian distributions, or deterministic mapping to the parameters of a categorical distribution [17]), which are not suitable for modeling non-Gaussian and heteroscedastic real-world medical data. In addition, they are still susceptible to degradation when faced with covariate shifts, including previously unseen noisy images and adversarial attacks.
@@
This paper introduces LaDiNE, Latent-guided Diffusion Nested-Ensembles, a novel, robust, probabilistic ensemble learning model for medical image classification that incorporates both transformers and diffusion models, given their recent success in medical imaging contexts [20, 21]. LaDiNE is a parametric mixture model that encodes invariant and informative features as latent variables, and performs functional-form-free inference to estimate the predictive distribution. Specifically, transformer encoder blocks are used as a hierarchical feature extractor that learn invariant features from images for each mixture component. The diffusion models are used as flexible distribution estimators to estimate the component distributions conditioned on the invariant features. In our formulation, each mixture component is interpreted equivalently as a Bayesian network that encodes the dynamics of observed (e.g. images) and latent variables. LaDiNE is specifically designed to be: (i) Robust to covariate shift; (ii) Provide calibrated confidence estimates; (iii) Be resilient to gradient-based adversarial attacks.
@@
    This paper introduces LaDiNE, a novel ensemble learning method that encodes invariant features and estimates the predictive distribution as a mixture model without specific assumptions on its component functional form, enabling flexible distribution modeling with limited medical data and sufficient expressiveness to encode complex patterns.
@@
    Extensive evaluation of the method against many popular methods under covariate shifts, both through a detailed comparison of the results of the proposed method against single models (vertical comparison) and state-of-the-art ensemble methods (horizontal comparison). Specifically, the following covariate shift scenarios are examined: (i) images with Gaussian noise injection, (ii) images with lower resolution, (iii) images with lower color contrast, and (iv) images with adversarial perturbation. Empirical evidence shows that our method improves on existing methods, achieving higher classification accuracy than the baseline methods. Extensive ablation studies are provided in order to justify the design choices made in the paper.
@@ II Related Work
Transformers in Medical Imaging.  Although transformer-based models have not yet been extensively explored in the field of medical imaging (as compared to computer vision), they have enormous potential to improve the modeling of complex spatial relationships and variability present in clinical data [20]. Some recent papers have shown how incorporating transformers helps to improve prediction accuracy in various medical imaging tasks. Peiris et al[24] proposes a transformer-based architecture that can encode local and global spatial cues for 3D tumor segmentation and exhibits robust performance against the presence of image artifacts. Chen et al. [25] integrates multi-scale feature extraction into the transformer to achieve improved performance in image-based gastric cancer detection, with the ability to be robust to noise. Wang et al. [26] uses transformers to capture X-ray sinograms’ global characteristics and achieves enhanced performance against artifacts in sparse-view CT reconstruction. However, these methods do not explicitly accounts for covariate shifts and data noise in architecture design, thus may not be able to obtain the optimal performance when facing a significant distribution shift in input images at test time.
@@
Diffusion Models in Medical Imaging.  Diffusion models have recently emerged as powerful generative models for medical imaging applications due to their ability to generate high-quality images, while remaining robust to distribution shifts [21, 27]. While most of the current work is focused on medical image generation and reconstruction with diffusion models [28, 29, 30, 31, 32], some methods have been developed for medical image segmentation via generating the segmentation mask [33, 34]. In other work, Li et al. [35] uses frequency-domain filters to guide the diffusion model for structure-preserving image translation, achieving robust generalization capability. Kim et al. [36] incorporates diffusion models into a representation learning framework for vessel segmentation and shows superior results on noisy data. However, the benefits of developing diffusion models in the context of medical image classification have not been exploited yet, despite the power of their generalizability and robustness to a wide variety of complex distribution shifts at test time.
III The Proposed Method: LaDiNE
@@
This work focuses on establishing a robust and generalizable model for the context of medical image classification, where a model trained on “clean” images, would be required to be robust to substantial, unseen covariate shifts on input images. The proposed framework consists of several important components: (i) transformer encoders (TEs) derived from Vision Transformers (ViTs) [37], (ii) conditional diffusion models (CDM), and (iii) feed-forward networks (FFNs). In this section, we first give a high-level overview of the proposed method. Then we describe the notation of variables and the computational paths involved in each neural network. Finally, we introduce the proposed probabilistic model in Sec. III-A and the training procedure for those neural networks in Sec. III-B.
@@
Overview.  LaDiNE is an ensemble deep learning model specifically designed to be robust to covariate shifts, while providing high-quality prediction confidence. To this end, LaDiNE (i) leverages early transformer encoders and a mapping network in order to learn image representations that are robust across different environments and (ii) estimates component distributions through conditional diffusion models, free from fixed distributional assumptions (see Fig. 2 (a) for an illustration of the proposed method).
@@
Vision Transformer and Transformer Encoders.  We follow the architecture described in [37], and introduce some additional notation used here. The Vision Transformer (ViT) is the stack of $L$ transformer encoder (TE) blocks, where each TE block acts as a deterministic mapping from the input sequence $s_{\text{in}}\in\operatorname{\mathbb{R}}^{n\times d}$ to the output sequence $s_{\text{out}}\in\operatorname{\mathbb{R}}^{n\times d}$. We define $\text{TE}:\operatorname{\mathbb{R}}^{n\times d}\rightarrow\operatorname{\mathbb{R}}^{n\times d}$ to represent this mapping. When computing TE, the input image $\mathbf{x}\in\operatorname{\mathbb{R}}^{H\times W\times C}$ is first divided into patches. Each patch is flattened and projected to a lower-dimensional embedding space. Given the patch size $(P,P)$, the number of patches is $N=\frac{HW}{P^{2}}$. The flattened patches are then linearly transformed:
$\mathbf{x}_{p}=\text{Reshape}(\mathbf{x},(N,P^{2}\cdot C)),\ e_{0}=\mathbf{x}_{p}\mathbf{E},$ (1)
where $\mathbf{E}\in\operatorname{\mathbb{R}}^{P^{2}\cdot C\times d}$ is a learnable embedding matrix. For simplicity, we denote Eq. 1 as a one-step function $\text{Emb}(\cdot)$:
$e_{0}:=\text{Emb}(\mathbf{x}).$ (2)
To retain positional information, learnable position embeddings are added to the patch embeddings. In a conventional ViT, a class token $e_{0}^{\text{cls}}$ is prepended to the sequence of embedded patches $e_{0}$ to be used for classification tasks:
$e_{0}^{\text{wcls}}=\text{Concat}(e_{0}^{\text{cls}},e_{0})+\mathbf{E}_{\text{pos}},\quad\mathbf{E}_{\text{pos}}\in\operatorname{\mathbb{R}}^{(N+1)\times d},$ (3)
where $\mathbf{E}_{\text{pos}}$ is the position embedding.
The sequence $e_{0}^{\text{wcls}}$ is then passed through the stack of $L$ TE blocks, we have
$e_{k}^{\text{wcls}}=\text{TE}_{k}(e_{k-1}^{\text{wcls}}),\quad k=1,2,\dots,L,$ (4)
where $e_{k}^{\text{wcls}}$ is the output of the $k$-th TE block. Note an arbitrary $e_{k}^{\text{wcls}}$ ($k\geq 1$) is composed of two embeddings: $e_{k}^{\text{cls}}$ (the class token) and $e_{k}$ (the content). $e_{k}^{\text{cls}}$ is conditionally independent of $e_{k}$ given the embedding $e_{k-1}^{\text{wcls}}$. For simplicity, we define
$\text{Te}(e_{k}):=\text{TE}(\text{Concat}(e_{k}^{\text{cls}},e_{k}))=\text{TE}(e_{k}^{\text{wcls}}),$
$\text{given that}\ e_{k}\mathrel{\perp\!\!\!\perp}e_{k}^{\text{cls}}\mid e_{k-1}^{\text{wcls}}.$ (5)
The final hidden state corresponding to the class token is used for classification with a feed-forward network (FFN):
$\mathbf{y}_{\text{logit}}=\text{FFN}(e_{L}^{\text{cls}}).$ (6)
In  [38, 39], the authors presented an investigation of the invariance (to adversarial noise) and informative hidden layers in ViTs [38, 39]. Here, we explore the extent to which the representations of the early TE blocks are indeed robust to various image perturbations. To this end, a number of experiments are performed. Fig. 1 depicts the results where the L2 norms of the differences between clean input and noisy input representations are shown, under four types of covariate shifts. The results indicate a clear pattern where, under all conditions, early TE blocks learn more invariant features than the deeper blocks. This finding motivates the use of early TE blocks in the predictions in order to improve robustness performance. Specifically, here $\text{TE}_{k}$ is included where $k=1,2,\dotsc,K$ and $K<L$.
@@
III-A Probabilistic Predictive Model
We now show how proposed model can be represented as a graphical model (see Fig. 2 (b)).
@@ III-A Probabilistic Predictive Model
where $\Theta$ denotes all parameters in the mixture, $\pi_{k}$ is the mixture weight for the $k$-th component distribution $\mathbf{p}_{k}(\mathbf{y}|\mathbf{x})$ with $\sum_{k=1}^{K}\pi_{k}=1$ and $\pi_{k}\geq 0$. For completeness, the joint distribution of variables in each component according to the Bayesian network is factorized as follows:
$\mathbf{p}_{k}(\mathbf{x},e_{1:k},z_{k},\mathbf{y}):=p(\mathbf{y}$ $|z_{k},\mathbf{x})p(z_{k}|e_{k})$
$\prod_{i=2}^{k}p(e_{i}|e_{i-1})p(e_{1}|\mathbf{x})p(\mathbf{x}).$ (9)
The $k$-th predictive component can be further factorized as:
@@
where $\delta(\cdot)$ is the Dirac Delta function, and $\text{Emb}(\mathbf{x})$ is the embedding step as shown in Eq. 2 to produce $e_{0}$. Here, the conditional distribution $p_{\theta_{k}}(\mathbf{y}|z_{k},\mathbf{x})$ modeled by the DM is parameterized by $\theta_{k}$, and the mapping network $g_{\phi_{k}}(\cdot)$ is parameterized by $\phi_{k}$. The notation of the parameters in $\text{Emb}(\cdot)$ and $\text{Te}(\cdot)$ are omitted here as they are packed in the TE blocks’ parameters, which are estimated along the training of the ViT. After simplification of the $\delta$ distribution, the model becomes:

### v5

@@ Abstract
Once deployed, medical image analysis methods are often faced with unexpected image corruptions and noise perturbations. These unknown covariate shifts present significant challenges to deep learning based methods trained on “clean” images. This often results in unreliable predictions and poorly calibrated confidence, hence hindering clinical applicability. While recent methods have been developed to address specific issues such as confidence calibration or adversarial robustness, no single framework effectively tackles all these challenges simultaneously. To bridge this gap, we propose LaDiNE, a novel ensemble learning method combining the robustness of Vision Transformers with diffusion-based generative models for improved reliability in medical image classification. Specifically, transformer encoder blocks are used as hierarchical feature extractors that learn invariant features from images for each ensemble member, resulting in features that are robust to input perturbations. In addition, diffusion models are used as flexible density estimators to estimate member densities conditioned on the invariant features, leading to improved modeling of complex data distributions while retaining properly calibrated confidence. Extensive experiments on tuberculosis chest X-rays and melanoma skin cancer datasets demonstrate that LaDiNE achieves superior performance compared to a wide range of state-of-the-art methods by simultaneously improving prediction accuracy and confidence calibration under unseen noise, adversarial perturbations, and resolution degradation.
{IEEEkeywords}
@@ 1 Introduction
In the rapidly evolving domain of medical imaging analysis, deep learning has led to enormous advances in many clinical domains of interest [1, 2, 3, 4, 5, 6, 7, 8, 9], notably in tasks such as detection of diabetic retinopathy in eye fundus images [2], classification of skin cancer [1], and identification of cancerous regions in mammograms [7]. Despite the fact that recent methods have achieved unprecedented success in controlled experimental settings, their fundamental building blocks – deep neural networks (DNNs) – are known to be sensitive to slight distribution changes or covariate shifts, and vulnerable to attacks [10, 11, 12]. As a result, their application to real-world clinical contexts often results in significant performance degradation, including inaccurate predictions and poorly calibrated confidence estimates.
@@
Data augmentation is a widely used tool for improving the generalization and robustness of DNNs. In the field of medical imaging, however, where datasets are typically smaller than natural image datasets especially for rare diseases [13], conventional data augmentation strategies may not always be suitable and can even lead to degradation in the performance of DNNs [14]. Designing suitable augmentation strategies from small medical imaging datasets can be challenging, requiring careful design choices in order to incorporate suitable inductive biases that lead to robust and generalizable DNN architectures and learning algorithms.
@@
Real-world medical imaging applications, such as medical image classification, often face unpredictable and complex covariate shifts that are difficult to anticipate during training [15], leading to significant challenges in maintaining reliable performance of DNN-based methods. In the field of machine learning, ensemble methods are popular choices to enhance robustness and generalization, as they combine the predictions of multiple models, effectively reducing variance and mitigating overfitting [16]. By leveraging the strengths of diverse models, ensemble techniques such as bagging, boosting, and stacking can achieve better and more robust predictive performance over a single model [17]. In addition, deep ensemble methods have been shown to improve both predictive performance and the quality of uncertainty estimates by training multiple DNNs independently and averaging their predictions [18]. Other frameworks involve developing algorithms to select better ensemble members and distribute input-dependent weights to each member, aiming to reduce the effect of weak members and providing performance gains [19, 20]. Despite the potential robustness offered by ensemble methods, their application in medical image classification to address covariate shifts and confidence calibration remains limited. While in natural image classification, existing ensemble methods often rely on simplified component distribution assumptions, such as Gaussian distributions or deterministic mappings to the parameters of a categorical distribution [18], which may not adequately capture the complexity and heteroscedasticity of real-world medical imaging data.
@@
This paper introduces LaDiNE, Latent-guided Diffusion Nested-Ensembles, a robust ensemble learning model designed for mitigating the aforementioned challenges in medical image classification. LaDiNE is a parametric mixture model that incorporates transformers and diffusion models, leveraging their recent success in medical imaging contexts [21, 22]. It encodes invariant and informative features as latent variables, and performs functional-form-free inference to estimate the predictive distribution. Specifically, a subset of transformer encoder blocks of a Vision Transformer (ViT) is used as a hierarchical feature extractor that learn invariant features from images for each mixture component. The diffusion models are used as flexible density estimators to estimate the component densities conditioned on the invariant features. In our formulation, each mixture component is interpreted equivalently as a Bayesian network that encodes the dynamics of observed (e.g., images) and latent variables. LaDiNE is specifically designed to be: (i) Robust to covariate shift; (ii) Provide calibrated confidence estimates; (iii) Be resilient to gradient-based adversarial attacks.
@@
    This paper introduces LaDiNE, a novel ensemble learning method that encodes invariant features and estimates the predictive component distribution without specific assumptions regarding its component functional form using, for the first time, diffusion models. This additional component enables flexible and expressive modeling of complex distributions, especially important in the with limited medical data.
@@
    Extensive evaluations demonstrate that the proposed method outperforms a wide variety of state-of-the-art (SOTA) methods in terms of classification accuracy under diverse covariate shift scenarios, including (i) images with Gaussian noise injection, (ii) images with lower resolution, (iii) images with lower color contrast, and (iv) images with adversarial perturbation. Empirical evidence shows significant improvements in classification accuracy over baseline methods. Detailed ablation studies further justify the design choices made in the paper.
@@ 2 Related Work
Transformers in Medical Imaging.  While transformer-based models have been less extensively studied in medical image analysis compared to their widespread application in computer vision, they hold significant potential for capturing complex spatial relationships and the variability inherent in clinical data [21]. Some recent works have shown how incorporating transformers helps to improve prediction accuracy in various medical imaging tasks. Peiris et al[25] proposes a transformer-based architecture that can encode local and global spatial cues for 3D tumor segmentation and exhibits robust performance against the presence of image artifacts. Chen et al. [26] integrates multi-scale feature extraction into the transformer to achieve improved performance in image-based gastric cancer detection, with the ability to be robust to noise. Wang et al. [27] uses transformers to capture X-ray sinograms’ global characteristics and achieve enhanced performance against artifacts in sparse-view CT reconstruction. Almalik et al. [28] extends this direction by leveraging intermediate representations from ViT blocks to improve robustness against adversarial attacks. Although these transformer-based methods demonstrate some degree of robustness to specific forms of noise and artifacts, they do not explicitly address confidence calibration and uncertainty quantification under noisy conditions or distribution shifts in input images. As a result, their ability to assist in informed clinical decision-making under unknown covariate shifts remains limited.
@@
Diffusion Models in Medical Imaging.  Diffusion models have recently emerged as powerful generative models for medical imaging applications due to their ability to generate high-quality images, while remaining robust to distribution shifts [22, 29]. While most of the current work is focused on medical image generation and reconstruction with diffusion models [30, 31, 32, 33, 34], some methods have been developed for medical image segmentation via generating the segmentation mask [35, 36]. In other work, Li et al. [37] uses frequency-domain filters to guide the diffusion model for structure-preserving image translation, achieving robust generalization capability. Kim et al. [38] incorporates diffusion models into a representation learning framework for vessel segmentation and shows superior results on noisy data. In the field of machine learning, recent work by Clark and Jaini demonstrates that a text-to-image diffusion model can function as a zero-shot classifier [39]. Han et al. [40] further introduces a diffusion-based conditional generative framework tailored for classification and regression on natural images and tabular data, showing strong performance in uncertainty estimation. Despite their robustness and ability to provide instance-level predictive uncertainty [41, 40], which can aid in informed clinical decision-making, their application in medical image classification has yet to be explored.
3 The Proposed Method: LaDiNE
@@
This work focuses on establishing a robust and generalizable model for medical image classification, where a model trained on “clean” images would be required to be robust to substantial, unseen covariate shifts. The proposed framework consists of several important components: (i) transformer encoders (TEs) derived from one Vision Transformer (ViT) [42], (ii) conditional diffusion models (CDM), and (iii) feed-forward networks (FFNs). In this section, we first give a high-level overview of the proposed method. Then we describe the notation of variables and the computational paths involved in each neural network. Finally, we introduce the proposed probabilistic model in Sec. 3.1 and the training procedure for those neural networks in Sec. 3.2.
@@
Overview.  LaDiNE is an ensemble deep learning model specifically designed to be robust to covariate shifts, while providing high-quality prediction confidence. To this end, LaDiNE (i) leverages early transformer encoders from only one ViT and mapping networks in order to learn image representations that are robust across different environments and (ii) estimates component distributions through conditional diffusion models, free from fixed distributional assumptions (see Fig. 2 (a) for an illustration of the proposed method).
@@
Vision Transformer and Transformer Encoders.  We follow the architecture described in [42], and introduce some additional notation used here. The Vision Transformer (ViT) is the stack of $L$ transformer encoder (TE) blocks, where each TE block acts as a deterministic mapping from the input sequence to the output sequence. We define $\text{TE}:\operatorname{\mathbb{R}}^{n\times d}\rightarrow\operatorname{\mathbb{R}}^{n\times d}$ to represent this mapping. When computing TE, the input image $\mathbf{x}\in\operatorname{\mathbb{R}}^{H\times W\times C}$ is first divided into patches. Those patches are then flattened and projected into a lower-dimensional embedding space. This process, including the addition of positional encoding, can be simplified into a single embedding function $e_{0}=\text{Emb}(\mathbf{x})$. Before feeding into the first TE block, a class token (with its own positional encoding) is appended to the embedding $e_{0}$. For simplifying the whole process, we denote another function $\text{Te}(\cdot)$ as:
$\text{Te}(e_{k}):=\text{TE}(\{\text{class token},e_{k}\}),$ (1)
here $e_{k}$ is the embedding output (excluding the class token) generated by the $(k-1)$-th TE block. The class token generated by the last TE block is used for classification with a feed-forward network (FFN).
In Almalik et al. and Walmer et al., the authors presented an investigation of the invariance to adversarial noise [28] and informative hidden layers in ViTs [43]. Extending these findings, we investigate whether the representations extracted from the early TE blocks in ViTs remain robust under a range of image perturbations, beyond adversarial noise considered in prior work [28]. To this end, a number of experiments are performed. Fig. 1 depicts the results where the L2 norms of the differences between clean input and noisy input representations are shown, under four types of covariate shifts. The results indicate a clear pattern where, under all conditions, early TE blocks learn more invariant features than the deeper blocks. This finding motivates the use of early TE blocks in the predictions in order to improve robustness performance. Specifically, here $\text{TE}_{k}$ is included where $k=1,2,\dotsc,K$ and $K<L$.
@@
3.1 Probabilistic Predictive Model
The proposed method is visually depicted in Fig. 2 (b), which provides a graphical representation of the dependencies among variables. This representation enables the factorization of the predictive distribution into conditional components. In this section, we describe the parametrization of each factorized conditional distribution.
@@ 3.1 Probabilistic Predictive Model
where $\Theta$ denotes all parameters in the mixture, $\pi_{k}$ is the mixture weight for the $k$-th component distribution $\mathbf{p}_{k}(\mathbf{y}|\mathbf{x})$ with $\sum_{k=1}^{K}\pi_{k}=1$ and $\pi_{k}\geq 0$. The $k$-th predictive component can be further factorized as:
@@
where $\delta(\cdot)$ is the Dirac Delta function, and $\text{Emb}(\mathbf{x})$ is the embedding step to produce $e_{0}$. Here, the conditional distribution $p_{\theta_{k}}(\mathbf{y}|z_{k},\mathbf{x})$ modeled by the DM is parameterized by $\theta_{k}$, and the mapping network $g_{\phi_{k}}(\cdot)$ is parameterized by $\phi_{k}$. The notation of the parameters in $\text{Emb}(\cdot)$ and $\text{Te}(\cdot)$ are omitted here as they are packed in the TE blocks’ parameters, which are estimated along the training of the ViT. After simplification of the $\delta$ distribution, the model becomes: